OpenStudy (anonymous):

I am working on solving a systems of two linear equations in two variables using the substitution method... could someone help me on how to start this x+y=7 x-2y=-5

OpenStudy (anonymous):

Try to isolate x in x+y=7, then substitute in x-2y=-5.

OpenStudy (anonymous):

so would I start x=-y+7

OpenStudy (anonymous):

then take that and substitute it x(y+7)-2y=5

OpenStudy (anonymous):

yes. And then?

OpenStudy (anonymous):

or is it x(-y+7)-2y=5

OpenStudy (anonymous):

Actually, (x)-2y=-5 becomes (-y+7)-2y=5 right?

OpenStudy (anonymous):

so x goes away

OpenStudy (anonymous):

Because x=-y+7 you substitute x by (-y+7).

OpenStudy (anonymous):

Easy, isn't it?

OpenStudy (anonymous):

yes..... thank you. I have had such a tough time with this stuff. Thank you so much

OpenStudy (anonymous):

And then? What is the solution?

OpenStudy (anonymous):

ahmm do I multiply (-y+7) by -2y?

OpenStudy (anonymous):

You are almost there. (-y+7)-2y=-5 becomes -y+7-2y=5 there is no multiplication

OpenStudy (anonymous):

why did the -5 become a positve 5

OpenStudy (anonymous):

sorry, I forgot (-y+7)-2y=-5 becomes -y+7-2y=-5 there is no multiplication

OpenStudy (anonymous):

I think I have it... subtract -2 from 7

OpenStudy (anonymous):

ok.. thank you

OpenStudy (anonymous):

I understand it much better now... but that's how I get I get it and then forget it :)

OpenStudy (anonymous):

try again

OpenStudy (anonymous):

Only add or subtract like terms. -y+7-2y=-5 -y-2y +7=-5 -3y +7=-5

OpenStudy (anonymous):

?

OpenStudy (anonymous):

do you want me to solve these equations?

OpenStudy (anonymous):

I think you should watch some videos on Youtube. Just search for "Systems Using Substitution" there, and you will understand better this subject.

OpenStudy (anonymous):

ok.. thanks you... do they have the elimation method too.

OpenStudy (anonymous):

There is a bunch of videos there that will help you. Try it some time.

OpenStudy (anonymous):

I will right now. ..... I really appreciate your help. thank you again

OpenStudy (anonymous):

You're welcome.